OpenStudy (anonymous):

Write the equation of a line that is perpendicular to the given line and that passes through the given point. 4x – 12y = 2; (10, –1) need help with this problem please!!!

OpenStudy (anonymous):

can i get any help?

OpenStudy (anonymous):

slope = -A/B of orig line -4/-12 = 3 Slop of orig line = 3 slope of perp line = neg inverse of slope of orig line mperp = -1/3 now use pt slope form

OpenStudy (anonymous):

answer : y+1 = (-1/3)(x-10)